The mean life of a certain brand of auto batteries is 44 months with a standard deviation of 3 months. Assume that the lives of all auto batteries of this brand have ca bell-shaped distribution. Using the empirical rule, find the percentage of auto batteries of this brand that have a life of 41 to 47 months 38 to 50 months 35 to 53 months
1
Expert's answer
2021-10-25T16:13:04-0400

"\mu=44 \\

\sigma=3 \\

P(41<X<47) = P(X<47) -P(X<41) \\

= P(Z< \frac{47-44}{3}) -P(Z< \frac{41-44}{3}) \\

= P(Z< 1) -P(Z< -1) \\

= 0.8413 -0.1586 \\

=0.6827 \\

= 68.27 \; \% \\

\\

P(38<X<50) = P(X<50) -P(X<38) \\

= P(Z< \frac{50-44}{3}) -P(Z< \frac{38-44}{3}) \\

=P(Z< 2) -P(Z< -2) \\

= 0.9772 -0.0227 \\

= 0.9545 \\

= 95.45 \; \% \\

\\

P(35<X<53) = P(X<53) -P(X<35) \\

= P(Z < \frac{53-44}{3}) -P(Z< \frac{35-44}{3} ) \\

= P(Z< 3) -P(Z< -3) \\

= 0.9986 -0.0013 \\

= 0.9973 \\

= 99.73 \; \%"